OpenStudy (anonymous):

If the term of a loan is 7 ½ years, then how many months will you have to pay?

OpenStudy (anonymous):

7 years = 7 * 12 = 84 months 1/2 years = 1/2 * 12 = 6 months hence 71/2 years = 84 + 6 = 90 months Therefore you need to pay the loan for 90 months..

OpenStudy (anonymous):

or

OpenStudy (anonymous):

\[7 \frac{1}{2} = \frac{2 \times 7 +1}{2} = \frac{15}{2} \] \[\frac{15}{2} \times 12 = 15 \times 6 = 90 \] Therefore you need to pay the loan for 90 months

OpenStudy (anonymous):

@andreaardzz